OpenStudy (anonymous):

if y=sin^-1(5x), then dy/dx=

OpenStudy (yuki):

is it \[\sin^{-1}(5x)\] ?

OpenStudy (anonymous):

yea!

OpenStudy (anonymous):

is that the answer you came up with?

OpenStudy (yuki):

okay, so you understand that this is arcsin(5x) right ? if you know the formula, you can use that. what I recommend though, is implicit differentiation. are you familiar with it ?

OpenStudy (anonymous):

Erm...I have a pretty concise way of demonstrating the derivation of the derivative of arcsin, let me know if you'd like to see it ayanamarcell. :)

OpenStudy (anonymous):

implicit differeniation?..no im not familar with it.

OpenStudy (yuki):

you go like this y = arcsin(5x) so using inverse relations, sin(y) = 5x when you differentiate both sides, do you know what happens?

OpenStudy (anonymous):

Yuki, there's a better way! :D (Well, if he/she's not familiar with implicit differentiation that is)

OpenStudy (anonymous):

@Yuki: no not really

OpenStudy (anonymous):

Sorry Yuki, you were going there. :P But technically, you don't need implicit differentiation for this type of problem (check my file for clarification).

OpenStudy (yuki):

okay, I don't want to force you down another thing you need to learn, so let's see what QM tells us. I am very interested

OpenStudy (anonymous):

@QM: I cant open your doc

OpenStudy (anonymous):

Ahh. :/ Darn. Then let me put my derivation here:

OpenStudy (anonymous):

thanks Yuki

OpenStudy (anonymous):

\[y = \sin^{-1} x\]\[x = \sin(y)\]\[\frac{dx}{dy} = \cos(y)\] (invert to get dy/dx) \[\frac{dy}{dx} = \frac{1}{\cos(y)}\] Using the main pythagorean identity,\[\frac{dy}{dx} = \frac{1}{\sqrt{1 - \sin^2(y)}}\] And because sin^2(y) = x^2 from our initial conditions,\[\frac{dy}{dx} = \frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

Because you only have a 5x inside the operator instead of x, you can multiply everything by the derivative of 5x (chain rule) and you have your derivative!

OpenStudy (yuki):

Wow that was a GOOD one !

OpenStudy (anonymous):

(And, of course, the x^2 inside the square root will be a (5x)^2, to keep things consistent.) Thanks yuki!

OpenStudy (yuki):

ayanamarcell, the one with implicit differentiation will be very similar to this proof, so if you understand this one you should not have problem learning it.

OpenStudy (anonymous):

thanks for the help! that was really simple to understand.

OpenStudy (anonymous):

wat about sin^-1 sqr x